Terrific Tubers Potato Farm needs cash until it sells and delivers its potato crop. Spud National Bank extends a $300,000, 6% loan to Terrific Tubers on December 1, 2018, for which Terrific Tubers signs a note payable. The loan agreement states that all principal and interest will be repaid on March 31, 2019. Terrific Tuber’s fiscal year ends on December 31.

Select the journal entry to record the payment of the note payable on March 31, 2019.

a) Debit Cash $306,000

Debit Interest Expense $4500

Credit Interest Payable $10,500

Credit Notes Payable $300,000

b) Debit Notes Payable $300,000

Credit Cash $300,000

c) Debit Notes Payable $300,000

Debit Interest Payable $1500

Debit Interest Expense $4500

Credit Cash $306,000

d) Debit Notes Payable $300,000

Debit Interest Payable $1500

Credit Interest Expense $4500

Credit Cash $297,000
Answer #1

Answer

  • Notes issued on 1 Dec 2018 for $ 300000
  • Interest Payable as on 31 Dec 2018 = $ 300000 x 6% x 1/12 = $ 1500
  • Interest Expense from 1 Jan 2019 to 31 Mar 2019 = $ 300000 x 6% x 3/12 = $ 4500
  • Total Cash paid on maturity on Mar 31, 2019 = $ 300000 + 1500 + 4500 = $ 306,000
  • Correct Answer = Option ‘C’
    Debit Notes Payable $300,000

Debit Interest Payable $1500

Debit Interest Expense $4500

Credit Cash $306,000